### v3

# Sub-cycle Nanotip Field Emission of Electrons Driven by Air Plasma Generated THz Pulses

Abstract
Terahertz pulses generated by two-color laser plasmas have reported peak field strengths exceeding MV/cm, and when illuminating metal nanotips the near-field enhancement at the tip apex should result in extremely high bunch charges and electron energies via sub-cycle cold field emission. Here, electron emission from tungsten nanotips driven by THz pulses generated by a long filament air-plasma are reported. Electron energies up to 1.1 keV and bunch charges up to 2x electrons per pulse were detected, well below values expected for peak field calculated via the time averaged Poynting vector. Investigations revealed a failure in the use of the time-averaged Poynting vector when applied to long filament THz pulses, due to spatio-temporal restructuring of the THz pulse in the focus. Accounting for this restructuring significantly reduces the field strength to approximately 160 kV/cm, consistent with the observed electron bunch charges, peak energies and their dependence on the tip position in the THz focus. Despite these findings, our results surpass previous THz plasma-driven electron generation by an order of magnitude in both electron energy and bunch charge and a path to increasing these by an additional order of magnitude by modification of the THz optics is proposed.
The development of sources of high brightness femtosecond electron pulses have led to significant advances in ultrafast electron microscopy and diffraction, enabling time-resolved atomic-scale imaging of dynamic processes in materials science, chemistry, and biology. Filippetto et al. (2022); Arbouet, Caruso, and Houdellier (2018) The emission of electrons from metallic surfaces and structures is often achieved by single or multi-photon photoemission and sub-cycle field emission using near-infrared femtosecond laser pulses.Ropers et al. (2007); Schenk, Krüger, and Hommelhoff (2010); Bormann et al. (2010); Echternkamp et al. (2016) Early demonstrations of nanotip electron emission employed low pulse energy, high-repetition-rate near-infrared (NIR) pulses. Hommelhoff, Kealhofer, and Kasevich (2006); Hommelhoff et al. (2006) Recently, single-cycle pulses of THz frequency light with peak fields exceeding 100 kV/cm have been used for cold-field emission (CFE) directly from the Fermi surface of metal nanotips. Li and Jones (2016); Matte et al. (2021) The metal nanotips induce a lightning rod effect, resulting in a broadband field enhancement at the tip apex of the order where is the free space wavelength and is the tip radius. Houard et al. (2020); Behr and Raschke (2008); Podenok et al. (2006)
Sub-cycle CFE offers several advantages compared to over-the-barrier photoemission or Schottky emission, including higher peak currents, as it is not limited by material damage thresholds from photon-induced heating inherent in NIR photoemission.Wellershoff et al. (1999) Additionally, THz field emission directly accelerates electrons post-tunneling in the local field, reaching multiple keV in energy, unlike photoemission which requires acceleration through a high voltage biased region of many cm. During the flight time, the pulse broadens significantly in space and time due to space-charge effects.
Pioneering work by Herink et al. used air plasma generated THz pulses centered at 1 THz with peak fields of kV/cm, Herink, Wimmer, and Ropers (2014) which, when aided by a -40 V bias, produced electron energies of 120 eV and yielding electrons per bunch at 1 kHz from a 10 nm radius nanotip. Subsequently Li and Jones, using a 15 Hz, 20 mJ Ti:sapphire fs laser generating 450 kV/cm peak field THz pulses with a central frequency of 0.15 THz via tilted pulse front optical rectification in LiNbO3, achieved electron energies up to 5 keV.Li and Jones (2016) Most recently Matte et al. achieved bunch charges of electrons per pulse at 1 kHz repetition rate with peak THz fields of 298 kV/cm and a central frequency of 1.2 THz generated by tilted pulse front optical rectification in LiNbO3, reaching maximum energies of 3.5 keV.Matte et al. (2021) This electron count surpassed previous CFE results by 3 orders of magnitude and was found to be consistent with Fowler-Nordheim (FN) theory, verifying the extreme quasi-static nature of the light-matter interaction. Despite these impressive results, air-plasma sources, with reported field strengths in the MV/cm range,Stepanov, Hebling, and Kuhl (2003); Yeh et al. (2007); Hirori et al. (2011); Guiramand et al. (2022) hold the potential to increase both the energy and bunch charge by another order of magnitude. In this work, the capabilities of two-color air-plasma THz sources to emit high-energy electron bunches via sub-cycle field emission from metal nanotips are studied.
Sub-cycle tunneling of the electron in the extreme quasi-static interaction limit occurs at a rate given by the Fowler-Nordheim (FN) equation:Fowler and Nordheim (1928)
where and are FN constants, and respectively,Lange et al. (2020) is the local electric field, and is the field required to narrow the potential barrier to zero, approximately 14 GV/m for tungsten.Forbes (2008) Here, , accounts for image charges and exponentially suppresses the emission for .Forbes and Deane (2010)
To reach the extreme quasi-static limit for sub-cycle CFE, the local field enhancement provided by metallic nanotips typically requires apex radii below 100 nm. These nanotips benefit from significant field enhancement at the apex due to the lightning rod effect, where the applied electric field () is enhanced by redistributing electric charges to cancel the field inside the conductor.Thomas et al. (2015) This effect narrows the emission area to the tip’s apex, offering a distinct advantage over photoemission sources, which have emission areas as wide as the focused beam on the photocathode. This point-like emission area increases the electron beam’s brightness and coherence, critical for high-resolution electron imaging.Latychevskaia (2017)
In this study, the experimental apparatus, shown in Fig. 1 (a), made use of a Coherent Legend Elite Duo HE+ Ti:sapphire regenerative amplifier, which emits 5 mJ, 35 fs pulses at 1 kHz with a central wavelength of 795 nm. The NIR output was split with 90% driving THz generation and 10% to the probe line. A two-color laser plasma was generated by focusing the main beam with a 50 cm lens into a 100 m-thick Beta-Barium-Borate (BBO) crystal placed before the focus, facilitating second harmonic generation at 400 nm. A zero-order half-wave plate after the BBO aligned the fundamental and second harmonic beams to vertically polarized states, creating a plasma filament and emitting THz radiation.
The THz beam was collimated using a 4-inch off-axis parabolic mirror (OAPM) and focused by a 3-inch OAPM onto the tungsten nanotip. Detection techniques included direct beam visualization with an IRXCAM INO384 microbolometer array THz camera, pyroelectric power measurements using a Gentec-EO THZ5B-VANTA meter, and standard electro-optic (EO) sampling for waveform and frequency spectrum analysis using a m GaP crystal which was bonded to a mm inactive GaP substrate.Wu and Zhang (1997)
Electron emission experiments were conducted using a tungsten nanotip housed in a spherical cube vacuum chamber, maintained at Torr. The chamber featured optical access ports and electrical feed-throughs, with the nanotip mounted on a 3-axis translation stage for precise positioning. Matte et al. (2021) Tungsten nanotips were reliably produced with tip apex radii between 10-20 nm using electrochemical etching.Lucier (2004) A scanning electron microscope (SEM) image of nanotips used here is shown in Fig. 1 (b). Significant reshaping of the tips was observed during previous experiments,Matte et al. (2021) leading to an estimated radius during emission between 15-35 nm. This reshaping was suggested by observed emission instability immediately after installation of a fresh tip, with the electron emission being reduced by a factor of 2 over 24 hours. At pressures of Torr, this emission instability is likely due to the presence of N2 and the formation of tungsten nitride layer on the nanotip. High fields are known to cause field-assisted etching, leading to tip reshaping.Rezeq, Pitters, and Wolkow (2006)
A Kimball Physics model FC-73A Faraday cup detector with a 5 mm diameter aperture was placed 2 cm above the nanotip. The collection angle is then approximately 7 degrees, allowing us to detect only a subset of electrons emitted within this angle. The setup included suppression and retarding grids to filter secondary electrons and analyze the energy spectrum of the emitted electrons. Additionally, a direct electrical connection at the nanotip base allowed measurement of the total tunneling current using a Keithley 6517B electrometer. The Faraday cup permitted mesaurement of the energy spectrum of the electron beam by capturing only electrons with energy surpassing the applied retarding potential. By varying the bias voltage and observing changes in the detected current, an electron energy spectrum can be obtained. Applying a low bias voltage of -20 V to the suppression grid further reduced the detection of secondary and scattered electrons.
The THz emission efficiency and spatial focusing was optimized using a a pyroelectric power meter (Gentec-EO THZ5B-VANTA) and a microbolometer camera located at the THz focus. The time-domain profile of the detected THz pulse is shown in Fig. 2(a), with the Fourier spectra in the inset showing a spectrum peaked at 1.46 THz. The FWHM of the focused beam profile, extracted from Fig. 2(b), are m and m along the x- and y-axes, respectively, where the uncertainties were estimated using frequency-dependent beam waist variation and camera absorption profiles. As shown in Fig. 2(c), at a distance of 16 mm after the focus, the beam displays the expected doughnut shape observed in previous air-plasma experiments and attributed to a Bessel-Gauss mode of the pulse.Klarskov et al. (2013) Here, the observed asymmetry is attributed to astigmatism in the imaging system. Rasmussen et al. (2023)
To estimate the field strength using the pulse energy, we begin by considering the magnitude of the Poynting vector, which describes the rate at which electromagnetic energy flows through space and is given by , Griffiths (2017) where is the energy flux (power per unit area), is the permittivity of free space, is the speed of light, and is the electric field. The total electromagnetic energy is then expressed as:
For a Gaussian profile, the transverse integral simplifies to:
Assuming that the electric field can be written as , where represents a normalized time-domain trace of the electric field, the total energy can be expressed as:
Solving for , the peak electric field, we get:
Using a time-domain trace obtained from air biased coherent detection (ABCD), the envelope of the field is obtained through a Hilbert transform, yielding . ABCD is preferred for this measurement as it avoids material-dependent limitations of electro-optic crystals, such as frequency-dependent phase mismatch, dispersion, and multiple reflections within the crystal, which can lead to temporal smearing and distortions in EOS Jepsen2011. The uncertainty in the time integral of the pulse envelope, , was determined by comparing the measured temporal pulse duration to that of a transform-limited Gaussian pulse, which gives a bandwidth-limited duration of . We measured the pulse energy as using a calibrated power meter, which introduces a uncertainty, corresponding to . Using these values, the peak electric field was calculated to be , with lower bound and upper bounds of and , accounting for uncertainty. These fields, when applied to our nanotip, would be expected to yield electron bunches of approximately e-/pulse with peak energies exceeding 6 keV, motivating the experiment presented here.
Following these initial measurements, unexpected changes were observed in the relative timing of the probe and THz pulses during the optimization of the EO signal and crystal scanning. This suggests the presence of several pulses arriving staggered in time, likely due to different emission times in the plasma or different path lengths traveled by each point-emitter in the plasma filament. This timing behavior has been fully described in a recent paper.Paulino, Colmey, and Cooke (2024) Here, further evidence of this effect is described and its impact on our observed electron emission. Most importantly to our electron emission experiments, this observation indicates that the field strength provided using the time-averaged Poynting vector above may be significantly overestimated, since the power meter cannot distinguish between a single large pulse or several smaller ones being focused to the optical axis at different points in time and space. As electron emission is an excellent probe of the quasi-static peak THz field, our data will serve as another method of field strength characterization with which to compare these estimates.
In our electron emission experiments, the nanotip position was scanned along the z-axis, revealing an extended, filament like focus, as shown in Fig. 3 (a). This observation follows our expectation of a Bessel-Gauss beam described elsewhere.Klarskov et al. (2013) In this figure the data is interpolated, with a total of 243 points for an average of 35 grid values per cross-section. During this scan a maximum of 1.5x electrons per pulse were measured via the replenishment current at the base, though at other times during the experiment as many as 2x were observed. In Fig. 3 (b), an electron emission line segment across X is shown, with Gaussian and Lorentzian fits superimposed. The data is better described by a Lorentzian distribution, consistent with previous observations of air-plasma beam waists.Klarskov et al. (2013)
The electron energy spectrum was obtained by applying retarding voltages to the Faraday cup. Measurements were taken with and without silicon wafers in the collimated THz beam path to modulate the incident field. As shown in Fig. 4(a), the maximum electron energies detected were 1100 eV, 700 eV, and 450 eV for 1, 2, and 3 silicon wafers, respectively, far below the 6 keV expected from a 600 kV/cm source. These measurements were obtained by iteratively querying our electrometer, with each current value representing an integrative measurement of charge over time. The current averages and their standard deviation were calculated across five iterations, resulting in the error bars shown.
Upon tunneling, electrons undergo ballistic acceleration and gain energy proportional to , diminishing to over a distance comparable to .Li and Jones (2016); Miller et al. (1996) Consequently, the energy gained by the electrons () is equivalent to the work done by the local electric field, , where is the field enhancement factor, and is the elementary charge of the electron. This relationship, however, implies a large dependence on for the final electron energies, something which Li et al. found to disagree with experimental observations.Li and Jones (2016) To explain these findings, Li et al. used the fact that , for a hemisphere-capped cylinder of length and radius in a static field, is proportional to . Podenok et al. (2006) This value can be further simplified to remove the length dependence, , as retardation effects will limit the effective length of the emitter to .Kang, Kim, and Park (2009) Plugging this into the earlier expression gives:
Giving us direct dependence of incident light field and wavelength on the electron energy, independent of . Using our electron energies and Eq. 6, the THz field strength may once again be estimated. As this is a proportionality relationship, it can be compared with existing electron energy vs field strength data to obtain an empirical version. Matte et al. performed a similar energy scan, using THz generated by OR in LiNbO3, and determined field strengths via EO sampling and as well as using the Poynting method, displaying excellent agreement. While we also attempted to calibrate the field strength of the plasma emitted THz pulses using EO sampling, the results of this calibration were always much too low to be considered realistic, again perhaps due to the spatio-temporal profile of the THz pulse mentioned above and the much higher bandwidth of the plasma THz pulses. In Fig. 4 (b), it can be seen that for the data of Matte et al., equation 6 is better approximated by replacing the factor of with a factor of . Using this relationship, we then estimate the plasma generated THz peak field strengths are 112.4, 71.5, 46.0 kV/cm, for 1, 2 and 3 wafers respectively. Accounting for the 70% transmission of each silicon wafer, the maximum field produced by air-plasma is approximately 160 kV/cm, significantly lower than the previous estimate based on THz power of 600 kV/cm. Given the magnitude of this discrepancy, it is unlikely to be caused by the aforementioned spatio-temporal profile of the THz pulse, suggesting an overestimation of field strength using the Poynting method.
Fig. 5 (a) shows the measured bunch charges for the THz field strengths investigated, and a comparison to the expectation of FN theory. Assuming an emission duration of 0.4 ps and an emission area of 10 nm2, a field enhancement factor of 6900 was estimated for a local field of 70 GV/m. This value is significantly higher than similar experiments which reported values between 800 and 3800. Li and Jones (2016); Matte et al. (2021); Vella et al. (2021). Potential explanations include variations in the work function due to surface contamination, which might alter the local field and geometry of the tip; deviations in emission area due to changing tip morphology; the presence of multiple emission sites on the tip surface; and space-charge effects, which could have influenced our observed electron energy distributions by causing additional broadening. However, these effects alone are unlikely to fully account for the observed discrepancies.
Another possible explanation is the presence of multiple pulses arriving staggered in time. To further understand this, the ballistic acceleration of electrons in the enhanced illuminating field was simulated. Assuming the near-field fall-off given by for a hyperboloid shaped tip,Li and Jones (2016); Miller et al. (1996) the dynamics of our highest energy electrons are modeled, emitted at the peak electric field, thus undergoing the maximum possible acceleration. In this purely 1D simulation, the electron’s position and velocity are iteratively computed and updated at each time step. The code calculates the electric field at the current position, based on spatial decay and the passing pulse, to determine the force acting on the electron. The simulated temporal evolution of the electron energy is shown in Fig. 5 (b). The field enhancement factor () importantly scales as .Kim et al. (2008) As our tip radius is not precisely known, a range of tip radii are given, with a range of enhancement factors then found. Our simulations show that, for tips between 15-35 nm in radii and assuming a incident field of 110 kV/cm estimated above, a field enhancement between to could have produced our maximum detected electron energies of keV. This range is closer to our expectation based on similar experiments. We note that although the estimate based on electron energy is directly related to the highest incident field on the tip and thus remains unaffected by the presence of multiple staggered pulses, the same cannot be said for our electron counts. In the previous FN-based estimation of , it was assumed that the tip is illuminated by a single pulse with a peak field of 110 kV/cm every 1 ms, given by repetition rate of our laser. However, if this assumption is incorrect—such as in the case where multiple pulses with weaker fields arrive more frequently, our calculated would be significantly reduced, bringing it in line with the simulated values and expectation.
In conclusion, sub-cycle cold field electron emission from a tungsten nanotip using a two-color laser plasma was studied, achieving electron energies of 1.1 keV and bunch charges up to electrons per pulse. While THz peak field strengths up to 600 kV/cm were estimated using the Poynting flux method, observations were consistent with a much lower field of 160 kV/cm based on the relationship between incident field and electron energy. This discrepancy was attributed to spatio-temporal focusing effects arising from collimation in the long plasma filaments using standard off-axis parabolic mirrors, producing a flying THz focus.Paulino, Colmey, and Cooke (2024)
We note that preliminary results using a plastic axicon as a collimating optic showed great promise, with a more than fourfold increase in peak THz field compared to the standard off-axis parabolic mirror, suggesting THz fields of over 600 kV/cm. By adopting this collimation scheme for future experiments, the current electron emission record of e-/pulse is expected to be surpassed, promising bunch charges up to e-/pulse and peak energies exceeding 6 keV. These advancements position plasma-driven THz sources as a promising option for ultrafast electron emission and possible THz-driven ultrafast electron microscopy using high-brightness, coherent electron beams from metal nanotips.
The authors wish to acknowledge financial support from the Natural Sciences and Engineering Research Council of Canada (NSERC), Fonds de Recherche du Québec-Nature et Technologies (FRQNT) and the Canadian Foundation for Innovation (CFI).
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Benjamin Colmey: Conceptualization (equal); Formal analysis (equal); Investigation (lead); Validation (equal); Writing – original draft (lead); Writing – review and editing (lead). Rodrigo T. Paulino: Conceptualization (equal); Formal analysis (equal); Investigation (equal); Validation (equal); Writing – review and editing (equal). Gaspard Beaufort: Investigation (supporting). David G. Cooke: Conceptualization (equal); Formal analysis (equal); Funding acquisition (lead); Project administration (lead); Validation (lead); Supervision (lead); Writing – review and editing (equal).
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.